LASER-INDUCED FLUORESCENCE STUDIES OF THE JET-COOLED CARBON DIOXIDE AND NI-TROUS OXIDE CATIONS

\underline{MOHAMMED A. GHARAIBEH} and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Laser-induced fluorescence (LIF) and single vibronic level (SVL) emission spectra of the Ã^{2}Π_{u} - X̃^{2}Π_{g} system of ^{12}CO_{2}^{+} and ^{13}CO_{2}^{+} and the Ã^{2}Σ^{+} - X̃^{2}Π system of N_{2}O^{+} have been observed. The cations were produced in a pulsed electric discharge jet with a precursor mixture of the corresponding neutral molecule in high pressure argon. The LIF bands of ^{12}CO_{2}^{+} and ^{13}CO_{2}^{+} were partially rotationally analyzed to obtain band origins which yielded an accurate measure of the excited state vibronic energy levels. The energy levels of both isotopologues were fitted with a Renner-Teller model that included spin-orbit coupling, Fermi resonance and anharmonic terms. SVL emission spectra were also recorded for the ^{13}CO_{2}^{+} ion and the ground state energy levels fitted using the same Renner-Teller model. Due to excited state predissociation, only the low lying vibrational levels have been observed in the LIF spectrum of N_{2}O^{+}. Nitrous oxide cation SVL emission spectra were collected and a complete vibronic analysis of the X̃^{2}Π electronic state has been carried out using the Renner-Teller model. Our experimental data provide rigorous tests for future ab initio potential energy surfaces of these important cations.